A stone is thrown horizontally at a speed of 86.3m/s from the top of a cliff 10.6m high.

What is the horizontal distance from the base of the cliff where the stone lands.

*hint* you will have to solve for time in the cortical direction friest before you can solve for the answer.
1
Expert's answer
2020-12-21T11:01:38-0500

Time of falling "t" can be found from the kinematic equation (if there is no vertical initial velocity component):



"h = \dfrac{gt^2}{2}"

where "h = 10.6m" is the height, and "g = 9.81m\/s^2" is the gravitational acceleration. Thus, obtain:



"t = \sqrt{\dfrac{2h}{g}} = \sqrt{\dfrac{2\cdot 10.6}{9.81}} \approx 1.47s"

In this time the stone covers the following horizontal distance:


"d = v_0t = 86.3m/s\cdot 1.47s \approx 126.9m"

Here "v_0 = 86.3m\/s" is the initial horizontal velocity of the stone.


Answer. 126.9 m.